# Probing thermonuclear bursts and X-ray reflection features in Aql X-1 during 2024 outburst

### v1

@@ 4.1 Reflection modeling of persistent emission
Using the best-fit spectral parameters for the nthcomp model, the optical depth of the corona is estimated to be $\tau_{e}$ $\sim$4.2 during NuSTAR obs-1. Similarly for NuSTAR obs-2, the optical depth is estimated to be $\sim$8.4. The seed photons are up-scattered by the hot electrons in the Compton cloud and gain energy (Sunyaev & Titarchuk, 1980, 1985). The relative energy gained by photons during inverse Compton scattering can be estimated by the Compton-y parameter ((Rybicki & Lightman, 1979)),
@@
The self-consistent reflection model, relxillCp, is utilized to describe the reflection feature. This model provides the inner disk radius of $\sim 4.3$ RISCO. For a rotating neutron star with dimensionless spin parameters $a^{\ast}$ = 0.259, $\rm R_{ISCO}$ is defined as van der Klis (2004),
@@
where, $\rm f_{ang}$ is anisotropy correction factor and $\eta$ is accretion efficiency, and $k_{A}$ is geometric coefficient. $M$ is the mass of the NS and $\text{F}_{bol}$ is the unabsorbed bolometric flux. We used 1–79 keV flux as bolometric flux, which is $\text{F}_{bol}\sim 5.2\times 10^{-9}$ erg cm-2 $s^{-1}$. The scaling factor $x$ can be estimated from $\text{R}_{\rm in}=\frac{xGM}{c^{2}}$. Assuming $f_{\rm ang}=1$, $\eta=0.1$, and $k_{A}=1$ (Cackett et al., 2009), the magnetic dipole moment of the NS can be estimated to be $\mu=9.6\times 10^{26}$ G cm3. The corresponding magnetic field strength at the poles of NS of radius 10 km is $B\leavevmode\nobreak\ \sim 1.9\times 10^{9}$ G. Similarly for the NuSTAR obs-2, the magnetic dipole moment and the magnetic field are estimated to be nearly $3.3\times 10^{26}$ G cm3 and $0.6\times 10^{9}$ G, respectively. The estimated magnetic field is consistent with the typical limit of other NS LMXBs (Mukherjee et al., 2015; Ludlam et al., 2016).
@@ 5 Summary and conclusions
We studied the broadband spectral and timing properties of Aql X-1 during its outburst in 2024 with NICER, Swift, and NuSTAR. We detected 6 type-I X-ray bursts with NICER and NuSTAR. The time-resolved spectra of bursts are modeled using an absorbed blackbody component, and the non-thermal persistent emission is modeled with the scaled background ($f_{a}$) approach. A soft excess was evident in the time-resolved burst spectroscopy, which indicates a deviation from the pure blackbody emission, possibly due to the enhanced mass accretion rate onto the NS. Alternatively, the soft excess can be modeled with the disk reflection model, and the contribution from the reflection model is found to be $\sim$20of the total emission as observed during the NICER burst. During the burst, the flux of the reflection component is found to be strongly correlated with the blackbody component flux. The bursts did not show any evidence of PRE. The preburst mass accretion rate is estimated to vary between $\sim$ (0.09-0.2) $\dot{M}_{Edd}$ for all six bursts. We also detected X-ray reflection features including a broad iron line at 6.4 keV and a Compton hump at around 20 keV with broadband NuSTAR, NICER, and Swift spectral study. The evolution of spectral parameters and disk properties are also investigated using the X-ray reflection spectral study at two different phases of the outburst. The accretion disk is found to be moderately ionized ($\log\xi\sim(2.6-3.2)$ erg cm s-1 )with a high disk density of $\log N\sim(18.6-19.5)$ cm-3. Assuming that the inner disk is truncated at the magnetosphere boundary, the magnetic field strength at the poles of the neutron star is estimated to be $\sim(0.6-1.9)\times 10^{9}$ G.

### v2

@@ 4.1 Reflection modeling of persistent emission
Using the best-fit spectral parameters for the nthcomp model, the optical depth of the corona is estimated to be $\tau_{e}$ $\sim$4.2 during NuSTAR obs-1. Similarly, for NuSTAR obs-2, the optical depth is estimated to be $\sim$8.4. The seed photons are up-scattered by the hot electrons in the Compton cloud and gain energy (Sunyaev & Titarchuk, 1980, 1985). The relative energy gained by photons during inverse Compton scattering can be estimated by the Compton-y parameter (Rybicki & Lightman, 1979),
@@
The self-consistent reflection model, relxillCp, describes the observed reflection feature. This model provides the inner disk radius of $\sim 4.3$ RISCO. For a rotating neutron star with dimensionless spin parameters $a^{\ast}$ = 0.259, $\rm R_{ISCO}$ is defined as van der Klis (2004),
@@
where, $\rm f_{ang}$ is anisotropy correction factor and $\eta$ is accretion efficiency, and $k_{A}$ is geometric coefficient. $M$ is the mass of the NS and $\text{F}_{bol}$ is the unabsorbed bolometric flux. We used 1–79 keV flux as bolometric flux, which is $\text{F}_{bol}\sim 5.2\times 10^{-9}$ erg cm-2 $s^{-1}$. The scaling factor $x$ can be estimated from $\text{R}_{\rm in}=\frac{xGM}{c^{2}}$. Assuming $f_{\rm ang}=1$, $\eta=0.1$, and $k_{A}=1$ (Cackett et al., 2009), the magnetic dipole moment of the NS can be estimated to be $\mu=9.6\times 10^{26}$ G cm3. The corresponding magnetic field strength at the poles of NS of radius 10 km is $B\leavevmode\nobreak\ \sim 1.9\times 10^{9}$ G. Similarly, for the NuSTAR obs-2, the magnetic dipole moment and the magnetic field are estimated to be nearly $3.3\times 10^{26}$ G cm3 and $0.6\times 10^{9}$ G, respectively. The estimated magnetic field is consistent with the typical limit of other NS LMXBs (Mukherjee et al., 2015; Ludlam et al., 2016).
@@ 5 Summary and conclusions
We studied the broadband spectral and timing properties of Aql X-1 during its outburst in 2024 with NICER, Swift, and NuSTAR. We detected 6 type-I X-ray bursts with NICER and NuSTAR. The time-resolved spectra of bursts are modeled using an absorbed blackbody component, and the non-thermal persistent emission is modeled with the scaled background ($f_{a}$) approach. A soft excess was evident in the time-resolved burst spectroscopy, which indicates a deviation from the pure blackbody emission, possibly due to the enhanced mass accretion rate onto the NS. Alternatively, the soft excess can be modeled with the disk reflection model, and the contribution from the reflection model is found to be $\sim$20of the total emission as observed during the NICER burst. During the burst, the flux of the reflection component is found to be strongly correlated with the blackbody component flux. The bursts did not show any evidence of PRE. The preburst mass accretion rate is estimated to vary between $\sim$ (0.09-0.2) $\dot{M}_{Edd}$ for all six bursts. We also detected X-ray reflection features, including a broad iron line at 6.4 keV and a Compton hump at around 20 keV with broadband NuSTAR, NICER, and Swift spectral study. The evolution of spectral parameters and disk properties are also investigated using the X-ray reflection spectral study at two different outburst phases. The accretion disk is found to be moderately ionized ($\log\xi\sim(2.6-3.2)$ erg cm s-1 ) with a high disk density of $\log N\sim(18.6-19.5)$ cm-3. Assuming that the inner disk is truncated at the magnetosphere boundary, the magnetic field strength at the poles of the neutron star is estimated to be $\sim(0.6-1.9)\times 10^{9}$ G.